OpenStudy (xxlyh):

can someone please check my answer? will medal! Which statement best explains whether the table represents a linear or nonlinear function? Input (x) Output (y) –2 7 –5 4 –8 1 –11 –2 (4 points) It is a linear function because there is a constant rate of change in both the input and output values. It is a nonlinear function because there is a constant rate of change in both the input and output values. It is a linear function because the input values are decreasing. It is a nonlinear function because the input values are increasing.

OpenStudy (xxlyh):

I chose D

OpenStudy (anonymous):

I am pretty sure this is correct..... idk for sure

OpenStudy (xxlyh):

@boldjon

OpenStudy (boldjon):

hold up

OpenStudy (boldjon):

A linear equation is, basically, a line. An example is y=2x+3. They are commonly found in the form y=mx+b where m is the slope, and b is the y-intercept (where the line crosses the y-axis). A non-linear equation can be many things. Parabolas and exponential equations are not linear

OpenStudy (boldjon):

If you had a table like this X | Y 1 | 2 3 | 4 5 | 3 a point is denoted as (x, y) use the slope formula with points (1,2) and (3,4) m = (4 - 2)/(3 - 1) = 2/2 = 1 thus we plug it into y=mx + b, we get y = 1*x + b now to find b we use (1,2) or we could use (3,4) 2 = 1*1 + b b = 2 - 1 b= 1 therefore the y intercept is 1 so we have the equation y = x + 1 now we test other points (5, 3), so we plug 5 into x to see if we get 3 out as a y value if we dont then the equation is non-linear y = 5 + 1 y = 6 thus the equation is not linear because 6 does not equal 3

OpenStudy (xxlyh):

I saw that, didn't help lol

OpenStudy (xxlyh):

it's either B or D

OpenStudy (boldjon):

i would go with b

OpenStudy (anonymous):

A function is linear if its graph forms a straight line. To know if the line would be straight, you have to examine the rate of change and whether or not it's constant. Every time you input a number, you add 9 to get the output. This means that the rate of change is constant- which, in turn, makes the system linear. So, the answer would be a: "It is a linear function because there is a constant rate of change in both the input and output values".

OpenStudy (xxlyh):

@ggw513 so it's not non linear?

OpenStudy (anonymous):

Yep

OpenStudy (xxlyh):

it was correct! I got 100% thaaaank you

OpenStudy (anonymous):

If you want to test that, you can turn the input and output into ordered pairs and graph it. For example, -2 and 7 would be (-2, 7) where -2 = x and 7 = y.

OpenStudy (anonymous):

No problem! Oh, and one more thing I forgot to mention- when it says that both the input and output have the same rate of change, it's referring to how as you continue down the table both sides keep decreasing by 3.

OpenStudy (anonymous):

Glad I could help :)